OpenStudy (madhatter):

David does landscaping for private homes and businesses. He charges $40.00 to travel to the site and then $40.00 per hour for his services. The equation y = 40 + 40x models the charge for his services y for any number of hours x. a. Graph the equation. Describe your graph. b. What is the slope of the line? What does the slope represent? c. What is the y-intercept of the line? What does the y-intercept represent? d. Julia paid David $240.00 for his services. How many hours did David work?

OpenStudy (madhatter):

plz help

jimthompson5910 (jim_thompson5910):

y = 40 + 40x y = 40 + 40(0) ... plug in x = 0 y = 40 + 0 y = 40 So one point on the graph is (0, 40)

jimthompson5910 (jim_thompson5910):

y = 40 + 40x y = 40 + 40(0) ... plug in x = 1 y = 40 + 40 y = 80 And another point on the graph is (1, 80)

jimthompson5910 (jim_thompson5910):

Plot the two points (0, 40) and (1, 80) Then draw a straight line through the two points to graph y = 40 + 40x

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (madhatter):

yep thanks can u help with the other parts too plz

jimthompson5910 (jim_thompson5910):

y = 40 + 40x is the same as y = 40x + 40

jimthompson5910 (jim_thompson5910):

y = 40x + 40 is in the form y = mx+b m = 40 (slope) b = 40 (y-intercept)

jimthompson5910 (jim_thompson5910):

What does the slope represent? it represents the idea that he's making $40 per hour (ie the total amount he's paid goes up $40 for each additional hour worked)

OpenStudy (madhatter):

oh ok thanks that makes alot of scene

jimthompson5910 (jim_thompson5910):

What does the y-intercept represent? The y-intercept is 40 as well (just a coincidence though), so this means that he started out with $40. This is the fixed amount he charges regardless of how many hours he worked

jimthompson5910 (jim_thompson5910):

part d) y = 40 + 40x 240 = 40 + 40x solve for x

OpenStudy (madhatter):

thanks so much this stuf dosent go in to my brain when i see the problem

OpenStudy (madhatter):

so the all the help u have given me is so nice thanks

jimthompson5910 (jim_thompson5910):

hmm well you gotta translate it to your own language somehow

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (madhatter):

ya lol